OpenStudy (anonymous):

To drive safely, the air pressure in your tires should be between 20 and 30 psi. Write this as an absolute value inequality, with x as the air pressure.

OpenStudy (anonymous):

here is a general way of solving such problems: |x-a|<=b -b<=x-a<=b in your case -b+a=20 b+a=30 2a=50 a=25 b=5 so it's |x-25|<5

OpenStudy (anonymous):

<= i meant

OpenStudy (anonymous):

Thanks.